OpenStudy (anonymous):

Using the equation P=m/V show that the equation PV=nRT can become P/p=(R/MW)T P=pressure p=density of air n=number of moles of gas R=gas constant T=temperature m=mass V=volume MW=mass of one mole of air

OpenStudy (mos1635):

PV=nRT PV=(m/MW)RT P(V/m)=(R/MW)T P(1/ρ)=(R/Mw)T Mw = 28.97 is the wean value if gas is air